OpenStudy (anonymous):

12x+4 = x+2 12x+4 is under a radical

OpenStudy (anonymous):

Sqroot(4*3*x+4*1)=x+2 2*Sqroot(3x+1)=x+2 2*Sqroot(3x+1)-x=2 x=8,0